A dish has 4 pennies, 2 nickels and 5 dimes. Randomly select 4 coins. What it the probability that the coins are all the same type? Show your work.(2 points)
Answer #1

Total~no.~of~coins=4+2+5=11\\\\ We~can~select~4~coins~from~11~coins~in~\binom{11}{4}~ways~and~if~all~4~coins\\ are~all~the~same~type~then~these~4~coins~are~either~pennies~and~dimes~so\\ total~no.~of~cases~that~the~4~coins~are~all~the~same~type\\ =\binom{4}{4}+\binom{5}{4}=1+5=6.\\\\ Required~probability=\frac{6}{\binom{11}{4}}=0.0182